OpenStudy (chazz16):

Look at the table of values below. x y 1 -1 2 -3 3 -5 4 -7 Which equation is represented by the table? A-y = 1 − 2x B-y = -x − 1 C-y = x − 2 D-y = 2x − 1

OpenStudy (eboogie4):

maybe A not sure

OpenStudy (texaschic101):

first take 2 points and find your slope using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (1,-1)...x1 = 1 and y1 = -1 (2,-3)...x2 = 2 and y2 = -3 now we sub slope(m) = (-3- (-1) / (2 - 1) slope(m) = (-3 + 1) / 1 slope(m) = -2/1 or just -2 now use y = mx + b and either of your points slope(m) = -2 (1,-1)..x = 1 and y = -1 and since we know the slope and we have our x and y points, we need to find the y intercept, the b position now we sub y = mx + b -1 = -2(1) + b -1 = -2 + b -1 + 2 = b 1 = b so your equation is : y = -2x + 1 or -y = 2x - 1

OpenStudy (mathmale):

I agree that it's best to derive an equation of this straight line from the given data, and then to compare your own equation to those listed as possible answers.